OpenStudy (moonlitfate):

Use trigonometric identities to transform the left side of the equation to the right side (0 < θ < π/2) : (tanθ)(cotθ) = 1

OpenStudy (noelgreco):

what is tan expressed in cos and sin?

OpenStudy (moonlitfate):

\[\tan \theta = \frac{ \sin \theta }{ \cos \theta}\] It makes a little more sense, but not completely.

OpenStudy (anonymous):

now what is cot expressed in cos and sin?

OpenStudy (anonymous):

in fact, you dont need to do this. do you know what cot is?

OpenStudy (anonymous):

\[\cot \theta = \frac{1}{\tan \theta}\]

OpenStudy (anonymous):

\[\tan \theta \times \cot \theta = \frac{\tan \theta}{\tan \theta} = 1\]